3. [OH] 2.3 x108 4. pOH 3.6
Answer #1

Criven [or] = 2.3x108 poh - -log (2.3x108) (:: poh - - 1oy Core] POH = -log(2-3) -10g (168) = -0.362 -(-8) log to = 8-0.362 =pOH = 36 Then we know that poH = -log [ok] > 10g [or] = -poh [04] = antilog (-poH) [ot] = 10 Pok :: [01] = 103.6 Cóm] = 0.25